OpenStudy (angelicahicks29):

(1-cos^2x)(1+cos^2x)=2sin^2x-sin^4x

OpenStudy (zehanz):

You have to prove the identity \((1-\cos^2x)(1+\cos^2x)=2\sin^2x-\sin^4x\). Maybe you have already tried multiplying out the LHS. This is tempting, because it has the form \((a-b)(a+b)\), which is equal to \(a^2-b^2\). Here, it would be \(1-\cos^4x\). This is not so nice, because now you have a fourth power of cos, while you only need sines. This will work: remember that famous identity: \(\sin^2x+\cos^2x=1\)? It can be used to convert from cos to sin: \(\cos^2x=1-\sin^2x\). Put this in place of the \(\cos^2x\) in the LHS, and you will arrive quickly at the RHS!

OpenStudy (zehanz):

Just try it. If you don't succeed, I can help you with it!